OpenStudy (yb1996):

Find the natural length of a heavy metal spring, given that the work done in stretching it from a length of 2ft to 3ft is one half the work done in stretching it from 3ft to 4ft.

OpenStudy (yb1996):

@freckles would you mind helping me out with this problem?

OpenStudy (irishboy123):

the energy stored in the spring is given by \(E = \frac{1}{2} k (x - x_o)^2\) where \(x_0\) is it's natural length the energy stored is the work done. from here it's just a maths problem. so "the work done in stretching it from a length of 2ft to 3ft " is it's concomitant increase in energy which is \(\frac{1}{2} k (3 - x_o)^2 - \frac{1}{2} k (2 - x_o)^2\) pattern match for stretching it from a length of 3ft to 4ft k's cancel, so looks like it's just algebra from there, ie solve for \(x_o\) :p

OpenStudy (yb1996):

Do you know is there is a way to solve this with integration?

OpenStudy (yb1996):

So according ti your way, would I set up the problem like this? \[\frac{ 1 }{ 2 }k(3-x _{0})^{2}-\frac{ 1 }{ 2 }k(2-x _{0})^{2} = \frac{ 1 }{ 2 }[\frac{ 1 }{ 2 }k(4-x _{0})^{2}-\frac{ 1 }{ 2 }k(3-x _{0})^{2}]\]

OpenStudy (yb1996):

@IrishBoy123

OpenStudy (irishboy123):

does that not work for you? i'd cancel the k's leaving one equation, 1 unknown, ie \(x_o\) bish bosh!

OpenStudy (yb1996):

I'm having difficulty understanding what you're trying to say. So would I have to set the equation that you gave me to zero and then solve for \[x _{0}\]?

OpenStudy (yb1996):

@IrishBoy123

OpenStudy (irishboy123):

\[\frac{ 1 }{ 2 }k(3-x _{0})^{2}-\frac{ 1 }{ 2 }k(2-x _{0})^{2} = \frac{ 1 }{ 2 }[\frac{ 1 }{ 2 }k(4-x _{0})^{2}-\frac{ 1 }{ 2 }k(3-x _{0})^{2}]\] \[(3-x _{0})^{2}-(2-x _{0})^{2} = \frac{ 1 }{ 2 }[(4-x _{0})^{2}- (3-x _{0})^{2}]\] that's a quadratic

OpenStudy (yb1996):

Ok, thank you!